AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Tue Apr 15 06:17:20 2025
Author
*
:
Subject
*
:
11.30 CRYRING vacuum and temperature - attachments 1-2 DSSSD bias & leakage current - attachment 3 c. 1uA reduction in DSSSD#1 and DSSSD#2 leakage currents cf. yesterday but remain high cf.June 2024 FEE64 temps OK - attachment 4 System wide checks - clock fail aida16 - attachments 5-6 aida16 global clock status 0x6 - non-critical failure, often resolved by a power cycle. ADC data item stats - attachment 7 DSSSD #3 n+n Ohmic side FEE64 rates return to c. normal cf. yesterday per FEE64 Rate spectra - attachment 8 per FEE64 1.8.L spectra - attachments 9-10 pulser peak width aida11 55 ch FWHM, aida09 75 ch FWHM only DSSSD #3 aida11/aida12 have expected peak widths per FEE64 1.8.W spectra - 20us FSR - attachments 11-13 WR timestamps OK - attachment 14 17.29 DSSSD bias & leakage current - attachment 15 FEE64 temps OK - attachment 16 WR timestamps OK - attachment 17 ADC data item stats - attachment 18 per FEE64 Rate spectra - attachment 19 17.35 All histograms & stats zero'd Start run Data file APR25/R5 alpha background EG&G Ortec 448 Pulser amplituide 90,000 normalise 5.00 x5 attenuator IN INT OSC frequency 2Hz t_r 100ns tau_d 50us (1000us setting does not work) polarity + ASIC settings 2025Apr08-13.45.30 slow comparator 0x64 aida03, aida04, aida09, aida10 LEC/EC range setting -> MEC per FEE64 Rate spectra - attachment 20 Merger & Tape Server - attachment 21-22 20.41 DAQ continues Data file APR25/R5_7 DSSSD bias & leakage current - attachment 23 FEE64 temps OK - attachment 24 ADC data item stats - attachment 25 Merger & Tape Server - attachment 26-27 per FEE64 Rate spectra - attachment 19
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown